# Sup-norm bounds for Jacobi cusp forms

### v2

@@ 2.5. Jacobi forms
(15) $\langle f_{1},f_{2}\rangle_{\mathrm{Pet}}=\frac{1}{\sqrt{4m}}\int\limits_{\mathcal{F}_{\Gamma_{0}}}\sum_{\mu=0}^{2m-1}\varphi_{\mu,1}(\tau)\overline{\varphi_{\mu,2}(\tau)}\,\eta^{k-\frac{1}{2}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta}{\eta^{2}},$
where we recall that $\mathcal{F}_{\Gamma_{0}}$ is a fundamental domain for the quotient space $X_{\Gamma_{0}}=\Gamma_{0}\backslash\mathbb{H}$.
2.6. Asymptotics for Bergman kernels
In this subsection, we recall asymptotics of Bergman kernels associated to tensor powers of holomorphic line bundles on compact complex Kähler manifolds, which are used in section 4 to derive bounds for theta functions.
Let $(M,\omega)$ be a compact complex Kähler manifold of dimension $n$ with positive closed $(1,1)$-form $\omega$. Let $\mathcal{L}$ be a positive hermitian holomorphic line bundle on $M$ and let $H^{0}(M,\mathcal{L}^{\otimes{m}})$ denote the vector space of global holomorphic sections its $m$-th tensor power $\mathcal{L}^{\otimes{m}}$ for $m\in\mathbb{Z}_{\geq 1}$. Let $|\,\cdot\,|_{\mathcal{L}^{\otimes{m}}}$ and $\langle\,\cdot\,,\,\cdot\,\rangle_{\mathrm{L}^{2},\,\mathcal{L}^{\otimes{m}}}$ denote the pointwise hermitian metric and the $\mathrm{L}^{2}$-inner product on $H^{0}(M,\mathcal{L}^{\otimes{m}})$, respectively.
Let $\{s_{j}\}$ denote an orthonormal basis of $H^{0}(M,\mathcal{L}^{\otimes{m}})$ with respect to the $\mathrm{L}^{2}$-inner product. For any $z\in M$, the function
$B_{\mathcal{L}^{\otimes{m}}}(z)\coloneqq\sum_{j}|s_{j}(z)|_{\mathcal{L}^{\otimes{m}}}^{2}$
is called the Bergman kernel associated to the line bundle $\mathcal{L}^{\otimes{m}}$. We note that the above definition is independent on the choice of an orthonormal basis of $H^{0}(M,\mathcal{L}^{\otimes{m}})$.
Let
(16) $\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)\coloneqq-\frac{i}{2\pi}\partial_{z}\partial_{\overline{z}}\log|s(z)|_{\mathcal{L}}^{2}$
denote the curvature form of the line bundle $\mathcal{L}$ at the point $z\in M$, where $s$ is any meromorphic section of $\mathcal{L}$. At any $z\in M$, there exists a coordinate chart around the point $z$ such that
$\omega(z)=\sum_{j=1}^{n}\frac{i}{2}\cdot\mathrm{d}z_{j}\wedge\mathrm{d}\overline{z}_{j}\quad\text{and}\quad\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)=\sum_{j=1}^{n}\frac{i}{2}\cdot\alpha_{j}\cdot\mathrm{d}z_{j}\wedge\mathrm{d}\overline{z}_{j}.$
The complex numbers $\alpha_{1},\ldots,\alpha_{n}$ are called the eigenvalues of the curvature form $\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)$ at the point $z\in M$. We set
$\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)\big{)}\coloneqq\prod_{j=1}^{n}\alpha_{j}.$
Since the line bundle $\mathcal{L}$ is positive, we have $\alpha_{j}>0$ for $j=1,\ldots,n$. Finally, we recall from Theorem 2.1 in [Bo96] the bound
(17) $B_{\mathcal{L}^{\otimes{m}}}(z)=\mathrm{det}_{\omega}\big{(}\mathrm{c}_{1}(\mathcal{L},|\,\cdot\,|_{\mathcal{L}})(z)\big{)}m^{n}+O(m^{n-1}),$
provided that $\mathcal{L}$ is a positive line bundle for any $z\in M$.
3. Sup-norm bounds for cusp forms revisited
@@
With notations as above, let $\Gamma\subset\mathrm{SL_{2}(\mathbb{R})}$ be a cofinite Fuchsian subgroup and $k\in\mathbb{R}_{\geq 5}$. Then, if $\Gamma$ is cocompact without elliptic elements, we have the bound
(29) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}(k).$
Moreover, if $\Gamma$ is cofinite, we have the bound
(30) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)}.$
The implied constants in the bounds (29) and (30) depend only on the Fuchsian subgroup $\Gamma$.
@@
When $\Gamma$ is cocompact without elliptic elements, the claimed bound (29) follows directly from Proposition 3.1.
@@
(31) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=\sup_{\begin{subarray}{c}\tau\in\partial\mathcal{F}_{Y}\\ Y=k/(2\pi)\end{subarray}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}},$
@@
With notations as above, let $\Gamma\subset\mathrm{SL_{2}(\mathbb{R})}$ be a cofinite Fuchsian subgroup. For $k\in\mathbb{R}_{\geq 5}$, let $f\in S_{k,\chi}(\Gamma)$ be a cusp form, which is normalized with respect to the Petersson inner product (9). If $\Gamma$ is cocompact without elliptic elements, we have the bound
(32) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma}(k).$
Moreover, if $\Gamma$ is cofinite, we have the bound
(33) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)}.$
The implied constants in the bounds (32) and (33) depend only on the Fuchsian subgroup $\Gamma$.
@@
With notations as above, let $\Gamma_{0}\subset\mathrm{SL_{2}}(\mathbb{R})$ be a fixed cofinite Fuchsian subgroup and let $\Gamma\subseteq\Gamma_{0}$ be a finite index subgroup of $\Gamma_{0}$. For $k\in\mathbb{R}_{\geq 5}$, let $f\in S_{k,\chi}(\Gamma)$ be a cusp form, which is normalized with respect to the Petersson inner product (9). If $\Gamma_{0}$ is cocompact without elliptic elements, we have the bound
(34) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma_{0}}(k).$
Moreover, if $\Gamma_{0}$ is cofinite, we have the bound
(35) $\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma_{0}}\big{(}k^{\frac{3}{2}}\big{)}.$
The implied constants in the bounds (34) and (35) depend only on the Fuchsian subgroup $\Gamma_{0}$.
@@
We are left to provide bounds for the pointwise Petersson norm of $f$, when $\tau$ ranges across the neighborhoods $\mathcal{F}^{Y}_{j}$ of the cusps. For this, we will have to distinguish between the two cases $Y>\frac{k}{2\pi}$ and $Y<\frac{k}{2\pi}$. Without loss of generality, we can assume that $j=1$, when $p_{1}\in\mathcal{P}_{\Gamma}$ is the cusp at infinity for $\Gamma$ lying above the cusp $p$ at infinity for $\Gamma_{0}$ with ramification index $[\Gamma_{0,p}:\Gamma_{p_{1}}]$ (note that $\Gamma_{0,p}$ denotes the stabilizer subgroup of $p$ in $\Gamma_{0}$). With the above notations, we obtain the inclusion $\Gamma\setminus\Gamma_{p_{1}}\subseteq\Gamma_{0}\setminus\Gamma_{0,p}$.
We first treat the case $Y>\frac{k}{2\pi}$, which implies that $\mathcal{F}_{1}^{Y}\subset\mathcal{F}_{1}^{k/(2\pi)}$. Arguing as in the proof of Theorem 6.1 in [FJK16], we deduce, recalling the inclusion $\Gamma\setminus\Gamma_{p_{1}}\subseteq\Gamma_{0}\setminus\Gamma_{0,p}$, that
$\quad\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\sup_{\tau\in\mathcal{F}_{1}^{k/(2\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
$\leq\frac{k-1}{4\pi}\Bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma\setminus\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\Bigg{)}$
$\leq\frac{k-1}{4\pi}\Bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\Bigg{)}.$
@@
$\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
$\quad\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r_{\Gamma_{0},k/(2\pi)}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma_{0},k/(2\pi)}/4)}\bigg{)}+\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}(m_{e_{j}}-1),$
@@
$r_{\Gamma_{0},k/(2\pi)}=\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)},\,\gamma\in\Gamma_{0}\setminus\bigg{(}\Gamma_{0,p}\cup\bigcup_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}\Gamma_{0,e_{j}}\bigg{)}\bigg{\}}>0.$
@@
$\frac{1}{\sinh^{2}(r_{\Gamma_{0},k/(2\pi)})}=O_{\Gamma_{0}}(1),$
@@
(39) $\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O_{\Gamma_{0}}(k).$
@@
(40) $\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O\big{(}k^{\frac{3}{2}}\big{)}.$
By means of (LABEL:thm6-eqn4), we thus deduce from (39) and (40) in the case $Y>\frac{k}{2\pi}$ the bound
@@
We finally turn to the case $Y<\frac{k}{2\pi}$, which implies that $\mathcal{F}_{1}^{k/(2\pi)}\subset\mathcal{F}_{1}^{Y}$. Here we find, arguing as in the preceding case that
$\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$ $\leq\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
$\leq\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
(42) $\quad\,+\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}.$
@@
$\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
@@
(44) $\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(2\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O\big{(}k^{\frac{3}{2}}\big{)}.$
By means of (42), we thus deduce from (43) and (44) in the case $Y<\frac{k}{2\pi}$ the bound
@@
If $f\in S_{k,\chi}(\Gamma)$ is not a Hecke eigenform, then there is no evidence from the literature to suggest that the estimates (32) and (33) can be improved. Thus the estimates (32) and (33) are expected to be optimal.

### v3

@@ 2.5. Jacobi forms
(15) $\langle\phi_{1},\phi_{2}\rangle_{\mathrm{Pet}}=\frac{1}{\sqrt{4m}}\int\limits_{\mathcal{F}_{\Gamma_{0}}}\sum_{\mu=0}^{2m-1}h_{\mu,1}(\tau)\overline{h_{\mu,2}(\tau)}\,\eta^{k-\frac{1}{2}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta}{\eta^{2}},$
where $\mathcal{F}_{\Gamma_{0}}$ denotes a fundamental domain for the quotient space $X_{\Gamma_{0}}=\Gamma_{0}\backslash\mathbb{H}$.
3. $L^{\infty}$-norm bounds for cusp forms revisited
@@
With notations as above, let $\Gamma\subset\mathrm{SL_{2}(\mathbb{R})}$ be a cofinite Fuchsian subgroup and $k\in\mathbb{R}_{\geq 5}$. Then, if $\Gamma$ is cocompact without elliptic elements, we have the $L^{\infty}$-norm bound
(27) $\|B_{k,\chi}\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}(k).$
Moreover, if $\Gamma$ is cofinite, we have the $L^{\infty}$-norm bound
(28) $\|B_{k,\chi}\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)}.$
The implied constants in the $L^{\infty}$-norm bounds (27) and (28) depend only on the Fuchsian subgroup $\Gamma$.
@@
When $\Gamma$ is cocompact without elliptic elements, the claimed $L^{\infty}$-norm bound (27) follows directly from Proposition 3.1.
@@
(29) $\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=\sup_{\begin{subarray}{c}\tau\in\partial\mathcal{F}_{Y}\\ Y=k/(4\pi)\end{subarray}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}},$
@@
With notations as above, let $\Gamma\subset\mathrm{SL_{2}(\mathbb{R})}$ be a cofinite Fuchsian subgroup. For $k\in\mathbb{R}_{\geq 5}$, let $f\in S_{k,\chi}(\Gamma)$ be an $L^{2}$-normalized cusp form. If $\Gamma$ is cocompact without elliptic elements, we have the $L^{\infty}$-norm bound
(30) $\|f\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma}(k).$
Moreover, if $\Gamma$ is cofinite, we have the $L^{\infty}$-norm bound
(31) $\|f\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)}.$
The implied constants in the $L^{\infty}$-norm bounds (30) and (31) depend only on the Fuchsian subgroup $\Gamma$.
@@
With notations as above, let $\Gamma_{0}\subset\mathrm{SL_{2}}(\mathbb{R})$ be a fixed cofinite Fuchsian subgroup and let $\Gamma\subseteq\Gamma_{0}$ be a finite index subgroup of $\Gamma_{0}$. For $k\in\mathbb{R}_{\geq 5}$, let $f\in S_{k,\chi}(\Gamma)$ be an $L^{2}$-normalized cusp form. If $\Gamma_{0}$ is cocompact without elliptic elements, we have the $L^{\infty}$-norm bound
(32) $\|f\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma_{0}}(k).$
Moreover, if $\Gamma_{0}$ is cofinite, we have the $L^{\infty}$-norm bound
(33) $\|f\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|f(\tau)\|_{\mathrm{Pet}}^{2}=O_{\Gamma_{0}}\big{(}k^{\frac{3}{2}}\big{)}.$
The implied constants in the $L^{\infty}$-norm bounds (32) and (33) depend only on the Fuchsian subgroup $\Gamma_{0}$.
@@
We are left to provide bounds for the pointwise Petersson norm of $f$, when $\tau$ ranges across the neighborhoods $\mathcal{F}^{Y}_{j}$ of the cusps. For this, we will have to distinguish between the two cases $Y>\frac{k}{4\pi}$ and $Y<\frac{k}{4\pi}$. Without loss of generality, we can assume that $j=1$, when $p_{1}\in\mathcal{P}_{\Gamma}$ is the cusp at infinity for $\Gamma$ lying above the cusp $p$ at infinity for $\Gamma_{0}$ with ramification index $[\Gamma_{0,p}:\Gamma_{p_{1}}]$ (note that $\Gamma_{0,p}$ denotes the stabilizer subgroup of $p$ in $\Gamma_{0}$). With the above notations, we obtain the inclusion $\Gamma\setminus\Gamma_{p_{1}}\subseteq\Gamma_{0}\setminus\Gamma_{0,p}$.
We first treat the case $Y>\frac{k}{4\pi}$, which implies that $\mathcal{F}_{1}^{Y}\subset\mathcal{F}_{1}^{k/(4\pi)}$. Arguing as in the proof of Theorem 6.1 in [FJK16], we deduce, recalling the inclusion $\Gamma\setminus\Gamma_{p_{1}}\subseteq\Gamma_{0}\setminus\Gamma_{0,p}$, that
$\quad\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\sup_{\tau\in\mathcal{F}_{1}^{k/(4\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}\leq\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
$\leq\frac{k-1}{4\pi}\Bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma\setminus\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\Bigg{)}$
$\leq\frac{k-1}{4\pi}\Bigg{(}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}+\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}\Bigg{)}.$
@@
$\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
$\quad\leq\frac{k-1}{2\pi}+\frac{3(k-1)}{\pi\cosh^{k-4}(r_{\Gamma_{0},k/(4\pi)}/4)}\bigg{(}1+\frac{1}{\sinh^{2}(r_{\Gamma_{0},k/(4\pi)}/4)}\bigg{)}+\frac{k-1}{4\pi}\sum_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}(m_{e_{j}}-1),$
@@
$r_{\Gamma_{0},k/(4\pi)}=\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)},\,\gamma\in\Gamma_{0}\setminus\bigg{(}\Gamma_{0,p}\cup\bigcup_{e_{j}\in\mathcal{E}_{\Gamma_{0}}}\Gamma_{0,e_{j}}\bigg{)}\bigg{\}}>0.$
@@
$\frac{1}{\sinh^{2}(r_{\Gamma_{0},k/(4\pi)})}=O_{\Gamma_{0}}(1),$
@@
(37) $\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O_{\Gamma_{0}}(k).$
@@
(38) $\frac{k-1}{4\pi}\sup_{\tau\in\partial\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O\big{(}k^{\frac{3}{2}}\big{)}.$
By means of (LABEL:thm6-eqn4), we thus deduce from (37) and (38) in the case $Y>\frac{k}{4\pi}$ the bound
@@
We finally turn to the case $Y<\frac{k}{4\pi}$, which implies that $\mathcal{F}_{1}^{k/(4\pi)}\subset\mathcal{F}_{1}^{Y}$. Again, arguing as in the proof of Theorem 6.1 in [FJK16], we deduce after recalling once again the inclusion $\Gamma\setminus\Gamma_{p_{1}}\subseteq\Gamma_{0}\setminus\Gamma_{0,p}$ that
$\sup_{\tau\in\mathcal{F}_{1}^{Y}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$ $\leq\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(4\pi)}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}$
$\leq\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
(40) $\quad\,+\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}.$
@@
$\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{0}\setminus\Gamma_{0,p}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}$
@@
(42) $\frac{k-1}{4\pi}\sup_{\tau\in\mathcal{F}_{1}^{Y}\setminus\mathcal{F}_{1}^{k/(4\pi)}}\sum_{\gamma\in\Gamma_{p_{1}}}\frac{1}{\cosh^{k}(\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)/2)}=O\big{(}k^{\frac{3}{2}}\big{)}.$
By means of (40), we thus deduce from (41) and (42) in the case $Y<\frac{k}{4\pi}$ the bound
@@
If $f\in S_{k,\chi}(\Gamma)$ is not a Hecke eigenform, then there is no evidence from the literature to suggest that the estimates (30) and (31) can be improved. Thus, the estimates (30) and (31) are expected to be optimal.